OpenStudy (anonymous):

Solve the equation 7+3x^(1/4) = 16

OpenStudy (akashdeepdeb):

Get 7 to RHS 3x^1/4 = 9 or (3x)^1/4 = (rt3)^1/4 CANCEL THE POWERS NOW!! get your answer!! :)

OpenStudy (anonymous):

i had 3x^1/4 = 9 why dont you divide 9 by 3

OpenStudy (akashdeepdeb):

Isn't the whole thing [(3x) raised to 1/4??]

OpenStudy (akashdeepdeb):

Okay if it is not then! You should get x^1/4 = 9/3 = 3 Now we know that 81^1/4 = 3 Put that in the equation! Thus x^1/4 = 81^1/4 Cut the powers!! GET X!! :)